# Premier-TACO is a Few-Shot Policy Learner: Pretraining Multitask Representation via Temporal Action-Driven Contrastive Loss

### v3

@@ B.2 Pretrained Visual Representations
PVR |
MVP |
R3M |
VC-1 |
PVR |
MVP |
R3M |
VC-1 |
Finger Spin |
Bin Picking |
Hopper Hop |
Disassemble |
Walker Walk |
Hand Insert |
Humanoid Walk |
Peg Insert Side |
Dog Trot |
Pick Out Of Hole |
Cup Catch |
Pick Place Wall |
Reacher Hard |
Shelf Place |
Cheetah Run |
Stick Pull |
Quadruped Walk |
Quadruped Run |
@@ B.3 LIBERO-10 success rate
| Unseen Tasks |
LfS |
Best PVRs |
Multitask |
Inverse |
CURL |
ATC |
SPR |
TACO |
Premier-TACO |
Task 0 |
Task 1 |
Task 2 |
Task 3 |
Task 4 |
Task 5 |
Task 6 |
Task 7 |
@@
Appendix C Implementation Details
Dataset For six pretraining tasks of Deepmind Control Suite, we train visual RL agents for individual tasks with DrQ-v2 Yarats et al. (2022) until convergence, and we store all the historical interaction steps in a separate buffer. Then, we sample 200 trajectories from the buffer for all tasks except for Humanoid Stand and Dog Walk. Since these two tasks are significantly harder, we use 1000 pretraining trajectories instead. Each episode in Deepmind Control Suite consists of 500 time steps. In terms of the randomly collected dataset, we sample trajectories by taking actions with each dimension independently sampled from a uniform distribution . For MetaWorld, we collect 1000 trajectories for each task, where each episode consists of 200 time steps. We add a Gaussian noise of standard deviation 0.3 to the provided scripted policy. For LIBERO, we take the human demonstration dataset from Liu et al. (2023), which contains 50 demosntration trajectories per task.
@@ Appendix C Implementation Details
For LIBERO, we use two randomly initialized (or pretrained) ResNet-18 encoders to encode the third-person view and first-person view images with FiLM (Perez et al., 2018) encoding method to incorporate the BERT embedding Devlin et al. (2019) of the task language instruction. During downstream behavior cloning, we apply a transformer decoder module with context length 10 on top of the ResNet encodings to extract the temporal information, and then attach a two-layer MLP with hidden size 1024 as the policy head. The architecture follows ResNet-T in Liu et al. (2023).
@@
Appendix D An Additional Ablation Study on Negative Example Sampling Strategy
@@
Appendix E Task instructions of downstream LIBERO tasks

### v4

@@ B.2 Pretrained Visual Representations
| PVR | MVP | R3M | VC-1 | PVR | MVP | R3M | VC-1 | ||
| Finger Spin | Bin Picking | ||||||||
| Hopper Hop | Disassemble | ||||||||
| Walker Walk | Hand Insert | ||||||||
| Humanoid Walk | Peg Insert Side | ||||||||
| Dog Trot | Pick Out Of Hole | ||||||||
| Cup Catch | Pick Place Wall | ||||||||
| Reacher Hard | Shelf Place | ||||||||
| Cheetah Run | Stick Pull | ||||||||
| Quadruped Walk | |||||||||
| Quadruped Run | |||||||||
@@ B.3 LIBERO-10 success rate
| Unseen Tasks | LfS | Best PVRs | Multitask | Inverse | CURL | ATC | SPR | TACO | Premier-TACO |
| Task 0 | |||||||||
| Task 1 | |||||||||
| Task 2 | |||||||||
| Task 3 | |||||||||
| Task 4 | |||||||||
| Task 5 | |||||||||
| Task 6 | |||||||||
| Task 7 | |||||||||
@@
Appendix C Additional Experiment Results on Downstream Online Reinforcement Learning
While our paper primarily focuses on the sample-efficient imitation learning for downstream adaptation, we can also apply reinforcement learning instead of imitation learning. Here in Figure 12, we have included additional experimental results on two unseen tasks, Walker Walk and Finger Spin, to showcase the downstream RL performance. We choose DrQ-v2 (Yarats et al., 2022) as the backbone visual RL algorithm and compare the performance of DrQ-v2 from scratch with DrQ-v2 using pretrained Premier-TACO encoder. Notably, representation learned from pretrained Premier-TACO encoder can also significantly accelerate downstream RL learning.
Appendix D Implementation Details
Dataset For six pretraining tasks of Deepmind Control Suite, we train visual RL agents for individual tasks with DrQ-v2 (Yarats et al., 2022) until convergence, and we store all the historical interaction steps in a separate buffer. Then, we sample 200 trajectories from the buffer for all tasks except for Humanoid Stand and Dog Walk. Since these two tasks are significantly harder, we use 1000 pretraining trajectories instead. Each episode in Deepmind Control Suite consists of 500 time steps. In terms of the randomly collected dataset, we sample trajectories by taking actions with each dimension independently sampled from a uniform distribution . For MetaWorld, we collect 1000 trajectories for each task, where each episode consists of 200 time steps. We add a Gaussian noise of standard deviation 0.3 to the provided scripted policy. For LIBERO, we take the human demonstration dataset from Liu et al. (2023), which contains 50 demosntration trajectories per task.
@@ Appendix D Implementation Details
For LIBERO, we use two randomly initialized (or pretrained) ResNet-18 encoders to encode the third-person view and first-person view images with FiLM (Perez et al., 2018) encoding method to incorporate the BERT embedding (Devlin et al., 2019) of the task language instruction. During downstream behavior cloning, we apply a transformer decoder module with context length 10 on top of the ResNet encodings to extract the temporal information, and then attach a two-layer MLP with hidden size 1024 as the policy head. The architecture follows ResNet-T in Liu et al. (2023).
@@
Appendix E An Additional Ablation Study on Negative Example Sampling Strategy
@@
Appendix F Task instructions of downstream LIBERO tasks